OpenStudy (anonymous):

sqrt(3)divided by sqrt(12)-sqrt(3) =

OpenStudy (saifoo.khan):

\[\frac{\sqrt3}{\sqrt12 - \sqrt3} \times \frac{ \sqrt 12 + \sqrt 3}{\sqrt 12 + \sqrt3}\]

OpenStudy (anonymous):

don't know how to get the answer though